# Position Paper: An Integrated Perspective on Data, Metrics, and Methodology for Deep Time-Series Forecasting

### v3

@@ B.1 Impact of Data Scale
B.2 Statistical and Gradient Boosting Decision Tree Baselines
To enhance the empirical robustness of our study, we integrate classical statistical models, including ARIMA (Makridakis & Hibon, 1997) and ETS (Hyndman & Athanasopoulos, 2018), along with the Gradient Boosting Decision Tree (GBDT) model, XGBoost, into the ProbTSframework. The results in Table 9 clearly demonstrate the superior performance of deep learning methods over simple statistical baselines, emphasizing the importance of capturing non-linear dependencies for accurate forecasts. Notably, ARIMA and ETS exhibit varied performance across different data characteristics. ARIMA struggles with datasets like Solar, characterized by weak trending and strong seasonality, while ETS shows better adaptability. Conversely, in cases of strong trending and weak seasonality, as observed in the ’Wikipedia’ dataset, ARIMA significantly outperforms ETS.
Utilizing the implementation from (Elsayed et al., 2021), we find that XGBoost competes well, even surpassing neural network models in certain scenarios. However, for datasets with more complex distributions like ’Solar’ and ’Electricity,’ advanced probabilistic estimation methods demonstrate a substantial advantage over traditional learning methods and point estimation techniques. This highlights the adaptability and strength of advanced probabilistic methods in handling intricate forecasting scenarios.
@@ B.2 Statistical and Gradient Boosting Decision Tree Baselines
Model Exchange Rate Solar Electricity Traffic Wikipedia
@@
B.3 Experiments on Univariate Datasets
In pursuit of a comprehensive analysis spanning univariate and multivariate scenarios, we examined a subset of M4 (Makridakis et al., 2020), M5 (Makridakis et al., 2022), and TOURISM datasets (Athanasopoulos et al., 2011)—crucial datasets for univariate time-series forecasting. Table 10 provides a quantitative assessment of the intrinsic characteristics of these new datasets, focusing on trending strength, seasonality, and data distribution complexity, as detailed in our paper. Notably, these datasets, except for M4-Daily may exhibit fewer seasonal patterns, do not introduce particularly unique characteristics.
@@ B.3 Experiments on Univariate Datasets
Table 11 presents experimental results for representative methods, consistent with our initial observations. Probabilistic estimation methods like GRU NVP and TimeGrad excel on datasets with complex distributions (e.g., M4-Weekly and M5), while simpler point forecasting methods such as DLinear and PatchTST perform well on datasets with relatively simple data distribution, like TOURISM-Monthly. Both autoregressive and non-autoregressive decoding schemes show comparable performance in short-term forecasting, as discussed in the main paper.”
@@
Model DLinear PatchTST GRU NVP TimeGrad
@@
B.4 Experiments on Synthetic Datasets
@@ B.6 Model Efficiency
Electricity 0.076 2.146 3.680 3.472 1.370
Traffic 0.078 2.149 15.926 8.298 1.390
Weather 0.075 2.145 3.085 0.574 1.721
Exchange 0.075 0.135 1.979 0.488 1.720
@@
Electricity 0.060 0.068 0.129 0.128 1.411
Traffic 0.161 0.168 0.361 0.333 9.102
Weather 0.004 0.012 0.021 0.012 0.070
Exchange 0.002 0.002 0.013 0.008 0.030
@@
Model pred len DLinear PatchTST LSTM NVP TimeGrad CSDI
ETTm1 96 0.0003 ± 0.0000 0.0003 ± 0.0000 0.0352 ± 0.0007 4.1067 ± 0.0504 16.3280 ± 0.0747
@@
B.7 Further Discussion on Cross-channel Interactions
@@
Customized neural architectures N-BEATS (Oreshkin et al., 2020)
N-HiTS (Challu et al., 2023)
Autoformer (Wu et al., 2021)
Informer (Zhou et al., 2021)
LTSF-Linear (Zeng et al., 2023) ✗/ ✓
PatchTST (Nie et al., 2023) ✗/ ✓
TimesNet (Wu et al., 2023)
Probabilistic estimation DeepAR (Salinas et al., 2020)
GP-copula (Salinas et al., 2019)
LSTM NVP (Rasul et al., 2021b)
LSTM MAF (Rasul et al., 2021b)
Trans MAF (Rasul et al., 2021b)
TimeGrad (Rasul et al., 2021a)
CSDI (Tashiro et al., 2021)
SPD (Bilos et al., 2023)

### v4

@@ D.1 Impact of Data Scale
D.2 Statistical and Gradient Boosting Decision Tree Baselines
To enhance the empirical robustness of our study, we integrate classical statistical models, including ARIMA [43] and ETS [31], along with the Gradient Boosting Decision Tree (GBDT) model, XGBoost, into the ProbTS framework. The results in Table 14 clearly demonstrate the superior performance of deep learning methods over simple statistical baselines, emphasizing the importance of capturing non-linear dependencies for accurate forecasts. Notably, ARIMA and ETS exhibit varied performance across different data characteristics. ARIMA struggles with datasets like Solar, characterized by weak trending and strong seasonality, while ETS shows better adaptability. Conversely, in cases of strong trending and weak seasonality, as observed in the ’Wikipedia’ dataset, ARIMA significantly outperforms ETS.
Utilizing the implementation from [22], we find that XGBoost competes well, even surpassing neural network models in certain scenarios. However, for datasets with more complex distributions like ’Solar’ and ’Electricity,’ advanced probabilistic estimation methods demonstrate a substantial advantage over traditional learning methods and point estimation techniques. This highlights the adaptability and strength of advanced probabilistic methods in handling intricate forecasting scenarios.
@@ D.2 Statistical and Gradient Boosting Decision Tree Baselines
Model Exchange-S Solar-S Electricity-S Traffic-S Wikipedia-S
@@
D.3 Experiments on Univariate Datasets
In pursuit of a comprehensive analysis spanning univariate and multivariate scenarios, we examined a subset of M4 [44], M5 [45], and TOURISM datasets [4]—crucial datasets for univariate time-series forecasting. Table 15 provides a quantitative assessment of the intrinsic characteristics of these new datasets, focusing on trending strength, seasonality, and data distribution complexity, as detailed in our paper. Notably, these datasets, except for M4-Daily may exhibit fewer seasonal patterns, do not introduce particularly unique characteristics.
@@ D.3 Experiments on Univariate Datasets
Table 16 presents experimental results for representative methods, consistent with our initial observations. Probabilistic estimation methods like GRU NVP and TimeGrad excel on datasets with complex distributions (e.g., M4-Weekly and M5), while simpler point forecasting methods such as DLinear and PatchTST perform well on datasets with relatively simple data distribution, like TOURISM-Monthly. Both autoregressive and non-autoregressive decoding schemes show comparable performance in short-term forecasting, as discussed in the main paper."
@@
Dataset DLinear PatchTST GRU NVP TimeGrad
@@
D.4 Experiments on Synthetic Datasets
@@ D.6 Model Efficiency
Electricity-L 0.076 2.146 3.680 3.472 1.370
Traffic-L 0.078 2.149 15.926 8.298 1.390
Weather-L 0.075 2.145 3.085 0.574 1.721
Exchange-L 0.075 0.135 1.979 0.488 1.720
@@
Electricity-L 0.060 0.068 0.129 0.128 1.411
Traffic-L 0.161 0.168 0.361 0.333 9.102
Weather-L 0.004 0.012 0.021 0.012 0.070
Exchange-L 0.002 0.002 0.013 0.008 0.030
@@
Dataset Pred len DLinear PatchTST LSTM NVP TimeGrad CSDI
ETTm1-L 96 0.0003 ± 0.0000 0.0003 ± 0.0000 0.0352 ± 0.0007 4.1067 ± 0.0504 16.3280 ± 0.0747
@@
D.7 Further Discussion on Cross-channel Interactions
@@
Customized neural architectures N-BEATS [52]
N-HiTS [12]
Autoformer [70]
Informer [74]
LTSF-Linear [72] ✗/ ✓
PatchTST [48] ✗/ ✓
TimesNet [69]
Probabilistic estimation DeepAR [58]
GP-copula [57]
LSTM NVP [56]
LSTM MAF [56]
Trans MAF [56]
TimeGrad [55]
CSDI [60]
SPD [8]